OpenStudy (anonymous):

What is x divided by 4/3 equals 6?

OpenStudy (oaktree):

8. Just do 4/3 times 6.

OpenStudy (oaktree):

You can multiply both sides by 4/3 and then you have x=6(4/3), which is 8.

OpenStudy (anonymous):

Okay, thanks so much!

OpenStudy (anonymous):

\[x/1*3/4=6\] \[3x/4=6\] \[3x=24\] \[x=\]

OpenStudy (anonymous):

8